OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____

OpenStudy (anonymous):

(x-8)^2

OpenStudy (anonymous):

by completing the square

OpenStudy (anonymous):

Fill in the missing term.

OpenStudy (anonymous):

take half of the middle term and square it

OpenStudy (anonymous):

So the answer is 16?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is half of 16?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

16/2

OpenStudy (anonymous):

16/2 is 4!

OpenStudy (anonymous):

its 8

OpenStudy (anonymous):

ohh yeah loll. i knew that.

OpenStudy (anonymous):

and then you square 8 to get the answer

OpenStudy (anonymous):

64

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

for trinomial ax^2+bx+c we have discriminant b^2-4ac here in given question a=1,b=-16 we need c ok for perfect square we need the discriminant so find the discriminant of above (-16)^2-4(1)c=0 so solve for c 4c=16^2 c=16^2/4 hope it is clear

OpenStudy (anonymous):

that how you complete the square

OpenStudy (anonymous):

take half the middle term and square

OpenStudy (anonymous):

assuming that the first term coefficient is 1

OpenStudy (anonymous):

4x2 + ___x + 49 What about this one?

OpenStudy (anonymous):

@MeganLewis close the current Question and ask new question . thanks!